Mr. A. is admitted to the medical unit with magnesium deficiency. His physician orders a magnesium...

  1. Mr. A. is admitted to the medical unit with magnesium deficiency. His physician orders a magnesium infusion. Order: Magnesium sulfate 4 g in 500 ml D5W at 500 mg/hr via infusion pump. What is the flow rate in ml/hr?
Answer #1

Given

Order:Magnesium sulphate

Dose:4g(4000mg)

Volume:500ml

Dose and time:500mg in 1 hour

If 500mg=1hour

Then 4000mg=??hour

Cross multiply and divide

=4000×1÷500

=4000÷500

=8 hours

Flow rate:Total volume (ml)÷Time(hr)

=500÷8

=62.5 or 63

Ans:63m/hr